Let f(x)=3−sqrtx^2+1f and g(x)=x−4.
Then (f∘g)(9)=    ,
(f∘g)(x)=
Answer #1

f(x)=3−sqrtx^2+1 = 3 - √(x2+1)

g(x) = x-4

(f∘g)(x) = f(g(x)) = f(x−4) = 3 - √((x-4)2+1) = 3 - √(x2-8x+17)

(f∘g)(9) = 3 - √(92-8(9)+17) = 3 -√(26) = 3 - 5.1 = -2.1